b) State whether the following statements are
TRUE or FALSE. Give reason(s) in support
of your answer.
i. Higher the marginal propensity to consume,
higher is the size of multiplier
ii. If investment is very sensitive to interest rate,
then we have a flat IS curve
(c) Use the following information (in rupees):
Income (Y) = 1,00,000
Nominal Money Supply (M) = 80,000
Price Level (P) = 20
Calculate the money growth rate required to
finance the budget deficit of Rs.10,000 in an
economy.
1
Expert's answer
2021-08-02T11:07:18-0400

b)i)true

The higher the marginal propensity to consume, the higher the multiplier,the more the increase in consumption from the increase in investment.

ii) true

If investment demand is highly sensitive to the interest rate, then a reduction in the interest rate causes a big increase in national income and product. Hence the IS curve is flat.

c) "Y=100,000"

"Money Supply( M)=80,000"

"Price Level (P)=20"

Money Growth Rate = "\frac {P\times Y}{M}"

"=\frac{20\times100,000}{80,000}"

"=25%" %